# ERA-IT: Aligning Semantic Models with Revealed Economic Preference for Real-Time and Explainable Patent Valuation

### v1

@@ 3.1 Sampling Strategy and Longitudinal Alignment
To establish a valuation framework with high external validity, it is critical to mitigate the selection bias often present in domain-specific studies. Unlike prior works restricted to high-velocity sectors, such as semiconductors or biotech, we aimed to capture the universal linguistic patterns of economic value applicable across the entire patent ecosystem. Accordingly, our data collection followed a rigorous, randomized, stratified sampling strategy. We targeted historical patent filings within the EPO database, specifically censoring the observation window to the 2010–2015 cohort. This timeframe ensures an ex post observation horizon of at least eight years, allowing for a complete retrospective analysis of economic survival trajectories. The final analytical corpus comprises 10,000 patent documents uniformly sampled across all primary IPC sections (A–H). By enforcing this heterogeneity, we ensure that the model learns to extract generalized economic signals rather than overfitting to the terminologies of specific industries. Table 1 presents representative examples derived from the constructed ground-truth dataset.
Table 1: Representative examples of the constructed ground-truth dataset. The corpus consists of 10,000 patents filed between 2010 and 2015, randomly sampled across diverse IPC sections (A-H). This randomized stratification ensures that the proposed framework captures universal value signals across heterogeneous technological landscapes, minimizing domain-specific overfitting.
@@
Patent 1 A (Human Necessities) 2011 Manual coffee brewing mechanism 2 yrs Class 1 (Low)
Patent 2 B (Transporting) 2013 Vehicle side mirror housing assembly 5 yrs Class 2 (Mid)
Patent 3 G (Physics/Computing) 2014 Neural network quantization method 9 yrs Class 3 (High)
Patent 4 C (Chemistry) 2012 CRISPR-Cas9 gene editing vector 11 yrs Class 3 (High)
Patent 5 H (Electricity) 2010 Legacy 3G communication protocol 3 yrs Class 1 (Low)
@@ 3.2 Operationalizing Value via Revealed Preferences
To construct objective ground-truth labels for supervised learning, we operationalized patent renewal records as a proxy for private economic value, grounded in the theory of revealed preferences kabore2012 ; harhoff2003 . Theoretically, the payment of annual renewal fees constitutes a "costly signal". It reflects the patent holder’s internal assessment that the expected future returns (or option value) of the asset exceed the cumulative and monotonically increasing maintenance expenses pakes1986 . In the context of information systems, this financial decision serves as a definitive binary signal at each time step, encoding the latent utility of the invention free from the noise of "cheap talk". Unlike citation metrics, which suffer from significant lag in accumulation and strategic manipulation, renewal behavior provides a real-time, forward-looking indicator of value retention. This allows us to convert discrete financial transactions into a robust target variable for algorithmic prediction kumar2024 .
@@ 3.3 Label Discretization and Value Stratification
To convert the continuous-time series of renewal events into a format suitable for classification, the corpus was stratified into three discrete value tiers. This discretization aligns with the typical hazard rate distribution of patent longevity observed in the econometric literature kumar2024 ; kabore2012 .
  • Class 1 (Low Value - Early Screening): Patents renewed one to three times. This tier represents the lower quartile of the value distribution, characterizing inventions where rights were relinquished during the initial screening phase. In economic terms, this rapid abandonment signals that the patent holder realized the technology’s lack of commercial viability or technical obsolescence shortly after the grant.
  • Class 2 (Medium Value - Standard Cycle): Patents renewed four to six times. These patents exhibit a standard technological lifespan, reflecting moderate economic relevance. They represent the median behavior of the dataset, where the asset generates sufficient utility to justify maintenance costs during the initial commercialization phase but lacks the strategic importance to warrant long-term protection against escalating fees.
  • Class 3 (High Value - Strategic Assets): Patents renewed seven or more times. This category captures the "long tail" of high-value IP. These patents were maintained for an extended period despite the EPO’s typically high fee structure. Continued renewal at this stage serves as a strong signal of sustained competitive advantage, strategic blocking value, or significant revenue generation capability.
@@ 4 Methodology: ERA-IT Framework
In this section, we propose the ERA-IT framework. Unlike traditional discriminative approaches that treat patent valuation as a black-box classification task, our framework adopts a generative paradigm to model the causal reasoning process underlying economic valuation explicitly. By aligning LLMs’ semantic processing capabilities with objective market signals, we aim to bridge the gap between high-dimensional technical specifications and their realized private value. Figure 3 illustrates the overall pipeline of the proposed architecture.
@@ 4.1 Problem Formulation: From Classification to Causal Reasoning
Traditional patent valuation is typically formulated as maximizing the conditional probability $P(y|X)$, where $y\in\mathcal{Y}$ represents the discrete value tier given the patent document $X$. However, this formulation suffers from interpretative opacity; it fails to explicate "why" a patent holds value, often relying on spurious correlations within the text.
@@ 4.2.1 Structural Schema for Cognitive Alignment
To ensure the generated reasoning aligns with domain experts’ cognitive processes, we define a four-step reasoning schema as follows. This schema forces the model to sequentially evaluate the structural components of patent value before reaching a conclusion.
@@ 4.3 Domain Adaptation via Instruction Tuning
We fine-tune a Student LLM (Llama-3-8B) using the constructed $\mathcal{D}_{CoT}$. To effectively inject domain-specific economic logic while preserving the model’s general linguistic capabilities, we adopt a parameter-efficient training strategy using LoRA (Low-Rank Adaptation).
@@ 4.3.1 Instruction-Formatted Input Representation
The raw patent document $X$ is transformed into a structured prompt in hierarchical tag format. The input sequence $I$ is defined as follows:
@@
where $\mathcal{I}_{task}$ denotes the role-playing instruction (e.g., "You are an expert in IP analyst"), forcing the model to adopt the persona of a valuation expert.
@@ 4.3.2 LoRA-based Adaptation
We freeze the pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ and update the learning parameters by introducing low-rank approximation matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$, where the rank is $r\ll\min(d,k)$. The weight update is formulated as:
$W=W_{0}+\Delta W=W_{0}+\frac{\alpha}{r}BA$ (3)
This approach allows for securing domain alignment by training less than 0.1% of the parameters, preventing catastrophic forgetting of general reasoning capabilities.
@@ 5.1.1 Datasets and Evaluation Metrics
We utilized the dataset of 10,000 EPO patents constructed in Section 4.2. The dataset was stratified and randomly split into training (8,000), validation (1,000), and testing (1,000) sets. Given the potential class imbalance in real-world scenarios, we report not only Accuracy but also Macro-F1 score and the Matthews Correlation Coefficient (MCC), which is considered a more robust metric for imbalanced classification tasks.
@@ 5.1.2 Baselines
  • Traditional machine learning: TF-IDF+SVM and TF-IDF+RF serve as fundamental benchmarks for text classification.
@@ 5.1.3 Implementation Details
The Student AI (Llama-3-8B) was fine-tuned using LoRA with rank $r=16$, $\alpha=32$, and dropout $0.05$. We used the AdamW optimizer with a cosine learning rate schedule (initial lr=2e4lr=2e-4) and a warm-up ratio of 0.03. All experiments were conducted on a single NVIDIA H100 GPU, ensuring reproducibility.
@@ 5.2 Performance Analysis (RQ1)
The ERA-IT framework achieves superior performance across all metrics, as shown in Table 2. Notably, it outperforms Longformer, suggesting that generative reasoning is more effective for valuation than merely processing long contexts. Furthermore, ERA-IT surpasses GPT-5-mini, demonstrating that parameter-efficient fine-tuning with economic signals yields better domain alignment than the general reasoning capability of closed-source giant models.
@@ 5.3 Ablation and Mechanism Analysis (RQ2)
Table 3: Ablation study on model components. The removal of each component leads to a performance drop, confirming their contributions to the framework.
@@
The most significant performance drop (-5.5%) occurs when the rationale generation ($r$) is removed. This confirms that the generated reasoning acts as a "cognitive scaffold", forcing the model to attend to relevant technical and legal details before making a prediction. Removing instruction masking also degrades performance, indicating that focusing the loss solely on the output helps the model learn the generation task more efficiently.
@@ 5.4.1 Performance Across Technology Domains
To verify the generalizability claimed in the Introduction, we analyzed the model’s accuracy across different IPC sections. As showm in Table 5, the model maintains high accuracy ($80\%$) in Physics (G) and Electricity (H), and shows robust performance even in Chemistry (C) and Mechanics (B), confirming its applicability across heterogeneous patent landscapes.
Figure 5: Prediction accuracy across major IPC sections. The model demonstrates consistent performance regardless of the technological domain.
@@ 5.5 Qualitative Case Study
6 Conclusion
6.1 Summary of Research Findings
In this study, we addressed the persistent challenges of information asymmetry and interpretative opacity in the patent market by proposing the ERA-IT framework. Moving beyond traditional bibliometric approaches that suffer from significant temporal latency, we theoretically redefined patent renewal history as a revealed economic preference. This enabled us to establish a novel supervised learning paradigm that leverages objective market signals to guide the semantic decoding of LLMs. Our empirical analysis, conducted on a dataset of 10,000 EPO patents randomly sampled across diverse technological domains, substantiates the efficacy of this approach. The ERA-IT framework not only achieved a significant performance gain over state-of-the-art discriminative baselines (Macro-F1 +7.2%) and zero-shot LLMs but also demonstrated robust generalizability across heterogeneous IPC sections ranging from Chemistry to Physics. Crucially, the model successfully reverse-engineers the Economic Chain-of-Thought, providing logical rationales that bridge the epistemic gap between obfuscated technical specifications and their realized market value.
@@ 6.2 Theoretical Implications
First, we introduce revealed economic preference as a novel supervision benchmark for the AI domain adaptation. This study points out that traditional citation-based bibliometric indicators suffer from systemic latency and cannot provide effective signals for real-time valuation. In contrast, the patent renewal decision is a market action taken by the right holder based on private information, directly reflecting the patent’s expected economic value. The core theoretical contribution of this research lies in systematically demonstrating that patent renewal history—a discrete, observable market behavior—serves as an objective and timely ground truth for large language models in understanding the semantics of patent texts. This finding challenges the prevailing reliance on citation-based indicators, proving that financial maintenance decisions offer a more immediate and less noisy proxy for technological value. By directly mapping linguistic patterns to economic survival, our work transcends the label ambiguity inherent in traditional unsupervised text analysis. It provides a clear, stable anchor in economic logic for decoding complex technical documents, thereby expanding the theoretical understanding of how value is encoded within technical documentation. By algorithmically merging high-dimensional, unstructured technical text with low-dimensional yet information-rich economic behavior data, the ERA-IT framework improves the efficiency and precision with which firms convert technical information into strategic intelligence. This theoretically empowers firms to conduct context-adaptive innovation in an increasingly digital, intelligent, and dynamic environment.
@@
Second, we define the Economic Chain-of-Thought as the mechanism to bridge the cognitive gap between semantic representation and market value, extending the application of Generative AI from simple text summarization to Causal Reasoning in Economic Contexts. Our framework does not stop at the black-box prediction of "input text, output value." Its key theoretical innovation is proposing the concept of the Economic Chain-of-Thought and making it an explainable output that the model must generate. This concept theoretically aligns the semantic reasoning process of AI with the economic causal logic of human decision-makers. It requires the model not only to predict an outcome but also to reverse-engineer, from obscure legal-technical texts, the chain of reasoning that supports a specific economic decision (to renew or not). This essentially builds a transparent cognitive scaffold for the AI’s reasoning, making its outputs no longer opaque but rational arguments that decision-makers can review, question, and adopt. By showing that instruction tuning with domain-specific economic logic significantly outperforms standard pre-training, this study establishes a new benchmark for "Vertical AI" applications where domain expertise is scarce, high-dimensional, and requires logical transparency.
@@
Third, we establish a computable theoretical pathway for firms to achieve AI’s context-adaptive innovation. This study provides a concrete instantiation of this concept: by using a domain-specific, objective economic behavior signal (renewal) as the supervision signal for instruction tuning, it compels a general-purpose LLM to adapt its internal semantic representations to the logic of the specific context of IP valuation. This alignment is not superficial keyword matching but drives the model’s generative reasoning to conform to the behavioral logic of micro-economic agents, thereby achieving a cognitive-level adaptation of the AI system from "general-purpose" to "domain expert".
@@
Overall, our study builds a new paradigm to resolve a core contradiction in the field: the chasm between technical complexity and economic measurability.

### v2

@@ 4.1 Sampling Strategy and Longitudinal Alignment
To establish a valuation framework with high external validity, it is critical to mitigate the selection bias often present in domain-specific studies. Unlike prior works restricted to high-velocity sectors, such as semiconductors or biotech, we aimed to capture generalized linguistic patterns of economic value applicable across the entire patent ecosystem.
Accordingly, our data collection followed a rigorous, randomized, stratified sampling strategy. We targeted historical patent filings within the EPO database, specifically defining the observation cohort to the 2010–2015 period. This timeframe ensures an ex post observation horizon of at least eight years, allowing for a complete retrospective analysis of economic survival trajectories.
The final analytical corpus comprises 10,000 patent documents uniformly sampled across all primary IPC sections (A–H). By enforcing this heterogeneity, we ensure that the model learns to extract universal economic signals rather than overfitting to the technical terminologies of specific industries. Table 1 presents representative examples derived from the constructed ground-truth dataset.
Table 1: Representative examples of the constructed ground-truth dataset. The corpus consists of 10,000 patents filed between 2010 and 2015, randomly sampled across diverse IPC sections (A–H). This randomized stratification ensures that the proposed framework captures universal value signals across heterogeneous technological landscapes.
@@
Pat. 1 A (Human Necessities) 2011 Manual coffee brewing mechanism 2 yrs Class 1 (Low)
Pat. 2 B (Transporting) 2013 Vehicle side mirror housing 5 yrs Class 2 (Mid)
Pat. 3 G (Physics/Computing) 2014 Neural network quantization 9 yrs Class 3 (High)
Pat. 4 C (Chemistry) 2012 CRISPR-Cas9 gene editing vector 11 yrs Class 3 (High)
Pat. 5 H (Electricity) 2010 Legacy 3G communication protocol 3 yrs Class 1 (Low)
@@ 4.2 Operationalizing Value via Revealed Preferences
To construct objective ground-truth labels for supervised learning, we operationalized patent renewal records as a proxy for private economic value, grounded in the theory of revealed preferences (danish2020; harhoff2003; jou2018; porter2023). Theoretically, the payment of annual renewal fees constitutes a “costly signal.” It reflects the patent holder’s internal assessment that the expected future returns (or option value) of the asset exceed the cumulative and monotonically increasing maintenance expenses (pakes1986).
In the context of information systems, this financial decision serves as a definitive binary signal at each time step, encoding the latent utility of the invention free from the noise of “cheap talk.” Unlike citation metrics, which suffer from significant lag in accumulation and potential strategic manipulation, renewal behavior provides a real-time, forward-looking indicator of value retention. This enables converting discrete financial transactions into a robust target variable for algorithmic prediction (kumar2024).
@@ 4.3 Label Discretization and Value Stratification
To convert the continuous-time series of renewal events into a format suitable for supervised learning, we stratified the corpus into three discrete value tiers. This tri-modal discretization is not arbitrary; it is empirically grounded in the marginal cost-benefit analysis of the EPO’s progressive fee structure and the typical hazard rate distribution observed in our dataset (danish2020; jou2018; kumar2024; porter2023). By mapping renewal counts to specific economic regimes, we capture the latent utility of each asset as follows:
  • Class 1 (Low Value - Early Screening): Patents renewed one to three times. This tier represents the lower quartile of the value distribution. In this phase, maintenance fees are relatively nominal; thus, a decision to relinquish rights signals an early “fail-fast” realization. Economically, this suggests the patent holder identified a lack of commercial viability or technical obsolescence almost immediately following the grant, opting to avoid further sunk costs.
  • Class 2 (Medium Value - Standard Cycle): Patents renewed four to six times. These patents represent the corpus’s median behavioral cluster. They demonstrate sufficient utility to justify maintenance during the primary commercialization window. However, they are abandoned before reaching the seventh year, a critical inflection point at which EPO renewal fees begin to escalate sharply. This reflects “standard” assets that provide moderate protection but lack the strategic necessity to warrant high-cost long-term retention.
  • Class 3 (High Value - Strategic Assets): Patents renewed seven or more times. This category captures the long tail of high-value IP and top-tier assets. Given the exponentially increasing cost of maintaining EPO patents as they approach their second decade, continued renewal beyond the seventh year signals a robust trajectory toward long-term retention. These documents represent technologies with sustained competitive advantages, high licensing potential, or significant strategic blocking value that far outweighs the cumulative financial burden of maintenance.
The resulting distribution provides a balanced target variable that reflects the revealed economic conviction of the patent holders, ensuring the model learns to distinguish between transient technical disclosures and sustained commercial assets.
@@ 5 Methodology: ERA-IT Framework
In this section, we propose the ERA-IT framework. Unlike traditional discriminative approaches that treat patent valuation as a black-box classification task, our framework adopts a generative paradigm to explicitly model the causal reasoning process underlying economic valuation. By aligning LLMs’ semantic processing capabilities with objective market signals, we aim to bridge the gap between high-dimensional technical specifications and their realized private value. Figure 3 illustrates the overall pipeline of the proposed architecture.
@@ 5.1 Problem Formulation: From Classification to Causal Reasoning
Traditional patent valuation is typically formulated as maximizing the conditional probability $P(y|X)$, where $y\in\mathcal{Y}$ represents the discrete value tier given the patent document $X$. However, this formulation suffers from interpretative opacity; it fails to explicate “why” a patent holds value, often relying on spurious correlations within the text.
@@ 5.2.1 Structural Schema for Cognitive Alignment
To ensure the generated reasoning aligns with domain experts’ cognitive processes, we define a four-step reasoning schema. This schema forces the model to sequentially evaluate the structural components of patent value before reaching a conclusion:
@@ 5.3 Domain Adaptation via Instruction Tuning
We fine-tune a Student LLM (e.g., Llama-3-8B) using the constructed $\mathcal{D}_{CoT}$. To effectively inject domain-specific economic logic while preserving the model’s general linguistic capabilities, we adopt a parameter-efficient training strategy using LoRA (Low-Rank Adaptation).
@@ 5.3.1 Instruction-Formatted Input Representation
The raw patent document $X$ is transformed into a structured prompt in a hierarchical tag format. The input sequence $I$ is defined as follows:
@@
where $\mathcal{I}_{task}$ denotes the role-playing instruction (e.g., “You are an expert IP analyst”), forcing the model to adopt the persona of a valuation expert, and $\oplus$ denotes string concatenation.
@@ 5.3.2 LoRA-based Adaptation
We freeze the pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ and update the learning parameters by introducing low-rank approximation matrices $A\in\mathbb{R}^{d\times r^{\prime}}$ and $B\in\mathbb{R}^{r^{\prime}\times k}$, where the rank is $r^{\prime}\ll\min(d,k)$. The weight update is formulated as:
$W=W_{0}+\Delta W=W_{0}+\frac{\alpha}{r^{\prime}}BA$ (3)
where $\alpha$ is a scaling factor. This approach allows for securing domain alignment by training less than 0.1% of the parameters, preventing catastrophic forgetting of general reasoning capabilities.
@@ 6.1.1 Datasets and Evaluation Metrics
We utilized the dataset of 10,000 EPO patents constructed in Section 5.2. The dataset was stratified and randomly split into training (8,000), validation (1,000), and testing (1,000) sets. Given the potential class imbalance in real-world scenarios, we report not only Accuracy but also the Macro-F1 score and the Matthews Correlation Coefficient (MCC), which is considered a more robust metric for imbalanced classification tasks.
@@ 6.1.2 Baselines
  • Traditional machine learning: TF-IDF+SVM and TF-IDF+Random Forest serve as fundamental benchmarks for text classification.
@@ 6.1.3 Implementation Details
The Student AI (Llama-3-8B) was fine-tuned using LoRA with rank $r=16$, $\alpha=32$, and dropout $0.05$. We used the AdamW optimizer with a cosine learning rate schedule (initial lr=2e4lr=2e-4) and a warm-up ratio of 0.03. All experiments were conducted on a single NVIDIA H100 GPU.
@@ 6.2 Performance Analysis (RQ1)
The ERA-IT framework achieves superior performance across all metrics, as shown in Table 2. Notably, it outperforms the traditional Random Forest model based on TF-IDF features by 36.3% in accuracy, 33.1% in Macro-F1, and 88.1% in MCC. It outperforms the discriminative pre-trained language model Longformer by 11.5% in accuracy, 8.7% in Macro-F1, and 27.4% in MCC, suggesting that generative reasoning is more effective for valuation than merely processing long contexts. Furthermore, it outperforms the zero-shot performance of the generative LLM GPT-5-mini by 9.6% in accuracy, 8.3% in Macro-F1, and 21.5% in MCC, demonstrating that parameter-efficient fine-tuning with economic signals yields better domain alignment than the general reasoning capability of closed-source giant models.
@@ 6.3 Ablation and Mechanism Analysis (RQ2)
Table 3: Ablation study on model components. Removing each component results in a performance drop, confirming its contribution to the framework.
@@
The most significant performance drop (-6.4%) occurs when the rationale generation ($r$) is removed. This confirms that the generated reasoning acts as a “cognitive scaffold,” forcing the model to attend to relevant technical and legal details before making a prediction. Removing instruction masking also degrades performance, indicating that focusing the loss solely on the output helps the model learn the generation task more efficiently.
@@ 6.4.1 Performance Across Technology Domains
To verify the generalizability claimed in the Introduction, we analyzed the model’s accuracy across different IPC sections. As shown in Figure 5, the model maintains high accuracy ($80\%$) in Physics (G) and Electricity (H), and shows robust performance even in Chemistry (C) and Mechanics (B), confirming its applicability across heterogeneous patent landscapes.
Figure 5: Prediction accuracy across major IPC sections. The model demonstrates consistent performance across technological domains.
@@ 6.5 Qualitative Case Study
7 Theoretical, Methodological, and Practical Implications
As noted in Section 2, existing patent-valuation approaches are constrained in providing economically grounded, explainable ex-ante assessments. Their limitations are threefold. First, they are largely ex-post, relying on outcome variables that materialize with substantial lag (fu2025; liu2024a). Second, metrics based on renewal behavior, composite indices, or citation structures are calibrated on long-term patterns and make limited use of the semantic content within individual patent texts (svensson2022; ponta2021; park2023). Third, these indicators aggregate value at the firm or field level and fail to provide stepwise economic rationales for a patent’s value at filing. To overcome these shortcomings, our study proposes a novel framework, ERA-IT, which integrates deep semantic analysis with interpretable reasoning for true ex-ante valuation. Compared with existing research, our study has significant theoretical, methodological, and practical implications.
@@ 7.1 Theoretical Implications
First, we introduce revealed economic preference as a novel supervision benchmark for AI domain adaptation. We elucidate that traditional citation-based bibliometric indicators suffer from systemic latency and fail to provide effective signals for real-time valuation. In contrast, the patent renewal decision is a market action taken by the right holder based on private information, directly reflecting the expected economic value of the patent. The core theoretical contribution of this research lies in systematically demonstrating that patent renewal history, a discrete and observable market behavior, serves as an objective and timely ground truth for Large Language Models to decode the semantics of patent texts. This finding challenges the prevailing reliance on bibliometrics and shows that financial maintenance decisions offer a more immediate, less noisy proxy for technological value. By directly mapping linguistic patterns to economic survival, our work transcends the label ambiguity inherent in traditional unsupervised text analysis. It provides a stable anchor in economic logic for decoding complex technical documents, thereby expanding the theoretical understanding of how value is encoded within technical documentation. By algorithmically merging high-dimensional and unstructured technical text with low-dimensional yet information-rich economic behavior data, the ERA-IT framework theoretically empowers firms to conduct context-adaptive innovation in an increasingly digital and dynamic environment.
@@
Second, we define the Economic Chain-of-Thought as the mechanism to bridge the cognitive gap between semantic representation and market value, extending the application of Generative AI from simple text summarization to Generative Reasoning in Economic Contexts. Our framework transcends the black-box paradigm of simply mapping input text to output value. Its key theoretical innovation is proposing the concept of the Economic Chain-of-Thought as an explainable output that the model must explicitly generate. This concept theoretically aligns the semantic reasoning process of AI with the economic logic of human decision-makers. It requires the model not only to predict an outcome but also to reverse-engineer the chain of reasoning that supports a specific economic decision, such as the choice to renew or abandon a patent, from obfuscated legal and technical texts. This process builds a transparent cognitive scaffold for AI reasoning and transforms outputs from opaque predictions into rational arguments that decision-makers can review and adopt. By showing that instruction tuning with domain-specific economic logic significantly outperforms standard pre-training, this study establishes a new benchmark for vertical AI applications where domain expertise is scarce, high-dimensional, and requires logical transparency.
@@
Third, we establish a computable theoretical pathway for firms to achieve AI-driven context-adaptive innovation. This study provides a concrete instantiation of this concept by using a domain-specific and objective economic behavior signal, specifically renewal data, as the supervision signal for instruction tuning. This approach compels a general-purpose LLM to adapt its internal semantic representations to the logic of the specific context of IP valuation. This alignment is not superficial keyword matching but rather drives the model’s generative reasoning to conform to the behavioral logic of microeconomic agents, thereby enabling cognitive-level adaptation of the AI system from a general-purpose learner to a domain expert.
@@
Overall, our study establishes a new paradigm to resolve a core contradiction in the field: the chasm between technical complexity and economic measurability.